# Galactica: A Scientific Large Language Model in Geoscience

### v1

@@ Abstract
Large language models (LLMs) have achieved huge success for their general knowledge and ability to solve a wide spectrum of tasks in natural language processing (NLP). Due to their impressive abilities, LLMs have shed light on potential inter-discipline applications to foster scientific discoveries of a specific domain by using artificial intelligence (AI for science, AI4S). In the meantime, utilizing NLP techniques in geoscience research and practice is wide and convoluted, contributing from knowledge extraction and document classification to question answering and knowledge discovery. In this work, we take the initial step to leverage LLM for science, through a rather straightforward approach. We try to specialize an open-sourced LLM into geoscience, by further pre-training the model with a vast amount of texts in geoscience, as well as supervised fine-tuning (SFT) the resulting model with our custom collected instruction tuning dataset. These efforts result in a model GeoGalactica consisting of 30 billion parameters. To our best knowledge, it is the largest language model for the geoscience domain. More specifically, GeoGalactica is from further pre-training of Galactica – a top-performing LLM trained with a large number of scientific documents. We train GeoGalactica over a geoscience-related text corpus containing 65 billion tokens curated from extensive data sources in the big science project Deep-time Digital Earth (DDE), preserving as the largest geoscience-specific text corpus. Then we fine-tune the model with 1 million pairs of instruction-tuning data consisting of questions that demand professional geoscience knowledge to answer. We validate GeoGalactica on various geoscience examinations and geoscience-related open-domain questions evaluated by a group of senior geoscientists. GeoGalactica demonstrates the state-of-the-art performance in a diverse range of NLP tasks in geoscience, as well as revealing the potential of using geoscience-related tools. In this technical report, we will illustrate in detail all aspects of GeoGalactica, including data collection, data cleaning, base model selection, pre-training, SFT, and evaluation. We open-source our data curation tools and the checkpoints of GeoGalactica during the first of pre-training in https://github.com/geobrain-ai/geogalactica.
@@ 1 Introduction
Leveraging the vast amount of resources of scientific literature, particularly the data resources collected for the Deep-time Digital Earth (DDE) project, we can create, organize, and manage a large and comprehensive geoscience dataset targeted for all stages in large language model training. In particular, we have introduced GAKG [2], Deep Literature111https://ddescholar.acemap.info/, GSO222https://gso.acemap.info/, and other platforms as carriers and repositories of geoscience text knowledge. These concerted efforts have not only allowed us to accumulate a comprehensive geoscience data archive but also have served as foundations for constructing an extensive instruction-tuning dataset for geoscience-related questions, GeoSignal-v2, which has been employed in supervised fine-tuning (SFT). In addition, we have developed and customized a series of data-cleaning tools that allow us to automatically convert various forms of raw data, such as PDF files, forms, equations, knowledge graphs, etc., into clean texts suited as training corpus for large language models. To our best knowledge, our collected corpus has become the largest geoscience dataset.
we have then successfully further pre-trained a language model with 30B parameters, with Galactica-30B [12] as its base model. The resulting model is thus named as GeoGalactica, empowering various academic tasks in the geoscience field. With its 30 billion parameters, this model represents the culmination of further pre-training and supervised fine-tuning, making it the largest language model dedicated to the geoscience domain. Our experimental findings demonstrate that, compared to models of equivalent scale, GeoGalactica exhibits exceptional performance on GeoBenchmark [15]. Regarding human evaluation, our model showcases impressive competence in geoscience-related tasks when compared with 5 general language models (ChatGPT333https://chat.openai.com/, Yiyan444https://yiyan.baidu.com/, Qianwen555https://qianwen.aliyun.com/, MOSS666https://moss.fastnlp.top/, ChatGLM777https://chatglm.cn/).
@@
In the subsequent sections of this technical report, we will provide a detailed description of the data collection and cleaning processes, base model selection, pre-training, supervised fine-tuning, and extensive evaluations in the creation of GeoGalactica. Additionally, we are committed to promoting open science by making our data curation tools and pre-training checkpoints available to the research community through our GitHub repositories888The list of related tools, data, and codes can be found in https://github.com/geobrain-ai/geogalactica.
@@
In addition to establishing the academic mega-model in geoscience, our goal is to contribute to a broader research community. Specifically, the experiences documented in this paper provide evidence for further community understanding of several open questions in the literature.
@@
A Toolchain for Data Cleaning: A high-quality training dataset is crucial for successfully training large language models. Therefore, our contribution to the community includes developing an efficient academic data preprocessing toolchain to construct a clean training corpus from PDF documents 999The toolchain is open-sourced on Github repos: https://github.com/Acemap/pdf_parser and https://github.com/davendw49/sciparser. In addition, an online demo of this toolchain can be found at https://sciparser.acemap.info/.
@@
Primitive Explorations to Use Tools: As for training GeoGalactica to use tools, we also construct a set of supervised data Geotools for training GeoGalactica to use tools. We also open-source the codes and data on Github.101010https://github.com/zthang/geotools
@@
Training Details and pre-training Checkpoints: We conducted model training on the accelerator hardware provided by the Advanced Computing East China Sub-center. We will describe in detail the pre-training and SFT processes in the remainder of this paper. In addition, we are releasing the training checkpoints during the first of the pre-training process on Hugging Face.111111https://huggingface.co/geobrain-ai/geogalactica
@@
Model and data analysis process: In building a domain-specific LLM, the model and the data should be effectively evaluated and analyzed. We provide a set of analysis and visualization methods for the SFT data and the weights of the GeoGalactica, open-sourced on Github.121212https://github.com/dbylynn/GeoGalactica_Analysis
@@ 3 Preliminary and Vocabulary
DDE. An International Union of Geological Sciences (IUGS) “Big Science” program that seeks to harmonize global deep-time Earth data with new protocols, platforms, and programs.
@@
DDE Scholar. DDE Scholar is a literature platform, aiming to construct a knowledge information system for geoscience scholars, which, step-by-step goes through knowledge arrangement, knowledge mining and knowledge discovery.
@@ 4 Data Collection and Cleaning
The training corpus of Galactica primarily consists of literature related to computer science and biochemistry rather than earth science. This suggests that Galactica may lack sufficient knowledge in the field of geoscience. To address this, we have collected approximately six million research papers specifically focused on earth science. These papers were carefully selected by professional experts in the field. Furthermore, we have expanded the GeoSignal dataset based on K2 to better support natural language processing tasks in earth science research. This expanded dataset was used for fine-tuning the model after further pre-training. In the following sections, we will provide a detailed explanation of how our dataset was constructed.
@@ 4.1 The Customized Pre-training dataset: GeoCoprus
During this process, we commenced our data collection in early 2020 by gathering a list of journals in geoscience from LetPub 131313https://letpub.com.cn/. We identified the corresponding publishers’ websites using the journal names and ISSN to collect open-access data through web page parsing. This process continued towards 2023 when we collaborated with experts in various sub-disciplines of geoscience, we collect paper from high-quality SCI journals in the mathematical geosciences, metamorphic petrology, geochronology,geomagnetism and paleomagnetism, geomorphology, tectonics, stratigraphy, hydrogeology, geophysical, geothermics, igneous and geochemistry, surficial geochemistry, geological mapping, sedimentological, petroleum geology, paleontology, paleogeography, and mineralogy. In total, we integrated a recommended list of 849 DDE-recommended journals(Appendix B shows the distribution of the collected papers in geoscience).
We employed the journal name list to search for journal information and their publishers’ websites. Through web page scraping, we collected HTML pages and subsequently conducted data parsing to extract metadata from papers. For open-access articles, we matched them with the parsed DOI and the corresponding PDF from Sci-Hub141414https://sci-hub.se/. If no PDF was available, we downloaded it based on the URL.
@@
In conclusion, we obtained a total of 6,611,063 papers. Our data collection system operated through a distributed information fusion mechanism, utilizing an 8-workstation k8s cluster. Data collection was conducted using Scrapy-Redis 151515https://github.com/rmax/scrapy-redis framework. Additionally, we implemented compression techniques for HTML data to address challenges related to large-scale data storage.
@@
We utilized data processing and enhancement tools based on DeepShovel and K2 during this process. With the help of Grobid [49] and pdffigure2 [50], we provided a comprehensive parsing solution for extracting text, images, tables, formulas, and other data types from research paper PDFs. This was further enhanced by DeepShovel for parsing tables and formulas, resulting in the development of the SciParser tool. We plan to open-source this tool and share it on GitHub.
@@
Currently, we are utilizing Grobid161616Footnote for Gribid. as our PDF parsing tool. Grobid can accurately extract text from PDF documents and perform structured extraction of articles. It provides an outline of the text, forming an XML structure that allows for restoring the original PDF layout. Additionally, Grobid enables precise localization of images, tables, formulas, and other data types. With the provided bounding boxes, we can obtain the corresponding images using the PyMuPDF tool 171717https://github.com/pymupdf/PyMuPDF. Further leveraging the OCR recognition integrated into DeepShovel [46], we can convert tables, formulas, and other elements into Markdown format. The parsing process is completed by writing all the parsed content into a markdown file for further use. Throughout the entire process, for tables, we utilize the DeepShovel PDF Table Parser 181818https://github.com/ShaoZhang0115/Table-Extraction-for-Geoscience-Literature. This tool ensures the completeness and accuracy of the table content while preserving the table structure, making it convenient to reconstruct tables using Markdown. As for formulas, we employ an improved version of Latex-OCR 191919https://github.com/lukas-blecher/LaTeX-OCR for the recognition, converting the parsing results into the string format. We open-source our PDF parsing solution on GitHub 202020https://github.com/Acemap/pdf_parser.
@@ 4.2.1 Domain General Natural Language Instruction
For the general instruction learning data, we have integrated four platforms constructed under the initiative of DDE, supported by Acemap, and reconstructed the data accordingly.
@@
This two platforms can be understood as collections of papers and datasets. Therefore, the Related Paper (with abstract) and Reference resolution of Deep Literature, as well as the Reference resolution of DDE DataExpo, serve as excellent datasets for establishing referential relationships.
Using the text processing tool mentioned earlier, we explicitly employ a multi-threaded Grobid to process all PDF papers and convert them into an intermediate XML format. Within the converted XML, we identify the bibl_id of in-text citations and then locate the corresponding reference paper titles in the XML’s reference section.
@@
For the article’s title, we used the Wikipedia library in Python 212121https://github.com/goldsmith/Wikipedia, which supports accessing sections of a Wikipedia page. Each section’s title, text, and sub-sections form a nested structure. By recursively traversing each page, we obtained a list of triplets comprising each section’s level, title, and paragraph. The triplets are structured as (level, title, paragraph), where the level indicates the depth of nesting, the title represents the section’s title, and the paragraph contains the corresponding text content.
@@
To extract the Entity mentioned in the article, we use the requests library and the BeautifulSoup 222222https://www.crummy.com/software/BeautifulSoup/ library to scrape the Wikipedia page directly. We retrieve the text from all tags labeled "p" and "ul" and treat them as paragraphs. Next, within these paragraph tags, we search for tags labeled "a" with a href attribute starting with "/wiki/". These represent the highlighted blue hyperlinked sections within the text. We collect these entities and output the tuple (paragraph, entities).
@@ 4.2.3 Self-Instruct
According to Alpaca [51] and Baize [52], using problem seeds to generate answer from ChatGPT 232323In the data curation process and experiments throughout this paper, we use the 2023 March version of ChatGPT and 2023 March version of GPT-4 unless otherwise specified is an appropriate way to build instruction tuning data. In geoscience scenarios, we generate 1000 questions per subject under the geoscience, and we put the problem seeds on GeoGalactica’s Github Repo.
@@
| DDE Scholar | Title (with Abstract) | (abstract; title) | 2,690,569 | |
@@
| DDE DataExpo | Title | (abstract; title) | 216,036 | |
@@ 5.3 Tool Learning
Action Input: "New York, Weather, 3M years"
• Example 2 Question: "What is the definition of plate tectonics?"
@@ 6.1.2 MMLU
Fortunately, we came across Galpaca-30B on Hugging Face 242424https://huggingface.co/GeorgiaTechResearchInstitute/galpaca-30b, which significantly reduced the carbon emissions from our finetuning experiments. This model utilized Alpaca’s instructions to learn from the dataset and was applied to SFT on Galactica-30B. Upon horizontal comparison, Galpaca-30B performed notably worse than the original Galactica and GeoGalactica in the majority of disciplines. This indicates that instruction learning in the general domain can significantly impact the performance of specialized domain models during practical evaluations.

### v2

@@ Abstract
Large language models (LLMs) have achieved huge success for their general knowledge and ability to solve a wide spectrum of tasks in natural language processing (NLP). Due to their impressive abilities, LLMs have shed light on potential inter-discipline applications to foster scientific discoveries of a specific domain by using artificial intelligence (AI for science, AI4S). In the meantime, utilizing NLP techniques in geoscience research and practice is wide and convoluted, contributing from knowledge extraction and document classification to question answering and knowledge discovery. In this work, we take the initial step to leverage LLM for science, through a rather straightforward approach. We try to specialize an open-sourced LLM into geoscience, by further pre-training the model with a vast amount of texts in geoscience, as well as supervised fine-tuning (SFT) the resulting model with our custom collected instruction tuning dataset. These efforts result in a model GeoGalactica consisting of 30 billion parameters. To our best knowledge, it is the largest language model for the geoscience domain. More specifically, GeoGalactica is from further pre-training of Galactica – a top-performing LLM trained with a large number of scientific documents. We train GeoGalactica over a geoscience-related scientific text corpus containing 65 billion tokens, preserving as the largest geoscience-specific text corpus. Then we fine-tune the model with 1 million pairs of instruction-tuning data consisting of questions that demand professional geoscience knowledge to answer. We validate GeoGalactica on various geoscience examinations and geoscience-related open-domain questions evaluated by a group of senior geoscientists. GeoGalactica demonstrates the state-of-the-art performance in a diverse range of NLP tasks in geoscience, as well as revealing the potential of using geoscience-related tools. In this technical report, we will illustrate in detail all aspects of GeoGalactica, including data collection, data cleaning, base model selection, pre-training, SFT, and evaluation. We open-source our data curation tools and the checkpoints of GeoGalactica during the first of pre-training in https://github.com/geobrain-ai/geogalactica§.
@@ 1 Introduction
Leveraging the vast amount of resources of scientific literature’s meta-data, particularly the data resources collected for the OpenAlex 111https://openalex.org/, Web of Science 222https://www.webofscience.com/wos/, Semantic Scholar 333https://www.semanticscholar.org/, and Acemap 444https://www.acemap.info, we can create, organize, and manage a large and comprehensive geoscience dataset targeted for all stages in large language model training. In particular, we have introduced GAKG [2], Deep Literature555https://idea.acemap.info/, GSO666https://gso.acemap.info/, and other platforms as carriers and repositories of geoscience text knowledge. These concerted efforts have not only allowed us to accumulate a comprehensive geoscience data archive but also have served as foundations for constructing an extensive instruction-tuning dataset for geoscience-related questions, GeoSignal-v2, which has been employed in supervised fine-tuning (SFT). In addition, we have developed and customized a series of data-cleaning tools that allow us to automatically convert various forms of raw data, such as PDF files, forms, equations, knowledge graphs, etc., into clean texts suited as training corpus for large language models. To our best knowledge, our collected corpus has become the largest geoscience dataset.
we have then successfully further pre-trained a language model with 30B parameters, with Galactica-30B [12] as its base model. The resulting model is thus named as GeoGalactica, empowering various academic tasks in the geoscience field. With its 30 billion parameters, this model represents the culmination of further pre-training and supervised fine-tuning, making it the largest language model dedicated to the geoscience domain. Our experimental findings demonstrate that, compared to models of equivalent scale, GeoGalactica exhibits exceptional performance on GeoBenchmark [15]. Regarding human evaluation, our model showcases impressive competence in geoscience-related tasks when compared with 5 general language models (ChatGPT777https://chat.openai.com/, Yiyan888https://yiyan.baidu.com/, Qianwen999https://qianwen.aliyun.com/, MOSS101010https://moss.fastnlp.top/, ChatGLM111111https://chatglm.cn/).
@@
In the subsequent sections of this technical report, we will provide a detailed description of the data collection and cleaning processes, base model selection, pre-training, supervised fine-tuning, and extensive evaluations in the creation of GeoGalactica. Additionally, we are committed to promoting open science by making our data curation tools and pre-training checkpoints available to the research community through our GitHub repositories121212The list of related tools, data, and codes can be found in https://github.com/geobrain-ai/geogalactica.
@@
In addition to establishing the academic mega-model in geoscience, our goal is to contribute to a broader research community. Specifically, the experiences documented in this paper provide evidence for further community understanding of several open questions in the literature. Warning: The model in this manuscript might produce hallucinations and reader discretion is recommended.
@@
A Toolchain for Data Cleaning: A high-quality training dataset is crucial for successfully training large language models. Therefore, our contribution to the community includes developing an efficient academic data preprocessing toolchain to construct a clean training corpus from PDF documents 131313The toolchain is open-sourced on Github repos: https://github.com/Acemap/pdf_parser and https://github.com/davendw49/sciparser. In addition, an online demo of this toolchain can be found at https://sciparser.acemap.info/.
@@
Primitive Explorations to Use Tools: As for training GeoGalactica to use tools, we also construct a set of supervised data Geotools for training GeoGalactica to use tools. We also open-source the codes and data on Github.141414https://github.com/zthang/geotools
@@
Training Details and pre-training Checkpoints: We conducted model training on the accelerator hardware provided by the Advanced Computing East China Sub-center. We will describe in detail the pre-training and SFT processes in the remainder of this paper. In addition, we are releasing the training checkpoints during the first of the pre-training process on Hugging Face.151515https://huggingface.co/geobrain-ai/geogalactica
@@
Model and data analysis process: In building a domain-specific LLM, the model and the data should be effectively evaluated and analyzed. We provide a set of analysis and visualization methods for the SFT data and the weights of the GeoGalactica, open-sourced on Github.161616https://github.com/dbylynn/GeoGalactica_Analysis
@@ 3 Preliminary and Vocabulary
Deep Literature. Deep Literature is a literature platform, aiming to construct a knowledge information system for geoscience scholars, which, step-by-step goes through knowledge arrangement, knowledge mining and knowledge discovery.
@@ 4 Data Collection and Cleaning
The training corpus of Galactica primarily consists of literature related to computer science and biochemistry rather than earth science. This suggests that Galactica may lack sufficient knowledge in the field of geoscience. To address this, we have collected approximately six million research documents specifically focused on earth science. These papers were carefully selected by professional experts in the field. Furthermore, we have expanded the GeoSignal dataset based on K2 to better support natural language processing tasks in earth science research. This expanded dataset was used for fine-tuning the model after further pre-training. In the following sections, we will provide a detailed explanation of how our dataset was constructed.
@@ 4.1 The Customized Pre-training dataset: GeoCoprus
During this process, we commenced our data collection in early 2020 by gathering a list of journals in geoscience from LetPub 171717https://letpub.com.cn/. We identified the corresponding publishers’ websites using the journal names and ISSN to collect open-access data through web page parsing. This process continued towards 2023 when we collaborated with experts in various sub-disciplines of geoscience, we collect paper from high-quality SCI journals in the mathematical geosciences, metamorphic petrology, geochronology,geomagnetism and paleomagnetism, geomorphology, tectonics, stratigraphy, hydrogeology, geophysical, geothermics, igneous and geochemistry, surficial geochemistry, geological mapping, sedimentological, petroleum geology, paleontology, paleogeography, and mineralogy. In total, we integrated a list of 849 geoscience-related journals(Appendix B shows the distribution of the collected papers in geoscience).
We employed the journal name list to search for journal information and their publishers’ websites. Through web page scraping, we collected HTML pages and subsequently conducted data parsing to extract metadata from papers. For open-access articles, we matched them with the parsed DOI and the corresponding PDF from Sci-Hub181818https://sci-hub.se/. If no PDF was available, we downloaded it based on the URL.
@@
In conclusion, we obtained a total of 5,980,293 papers. Our data collection system operated through a distributed information fusion mechanism, utilizing an 8-workstation k8s cluster. Data collection was conducted using Scrapy-Redis 191919https://github.com/rmax/scrapy-redis framework. Additionally, we implemented compression techniques for HTML data to address challenges related to large-scale data storage.
@@
Notice: All the data employed in this manuscript is derived from web pages that are publicly available, thereby introducing a potential bias in the reliability of the data. Users are cautioned to be mindful of potential hallucination problems that may occur when utilizing large-scale models. Furthermore, this paper adheres to all copyright concerns. Should any issues arise, stakeholders are encouraged to notify the authors.
@@
We utilized data processing and enhancement tools based on DeepShovel and K2 during this process. With the help of Grobid [49] and pdffigure2 [50], we provided a comprehensive parsing solution for extracting text, images, tables, formulas, and other data types from research documents. This was further enhanced by DeepShovel for parsing tables and formulas, resulting in the development of the SciParser tool. We plan to open-source this tool and share it on GitHub.
@@
Currently, we are utilizing Grobid as our PDF parsing tool. Grobid can accurately extract text from PDF documents and perform structured extraction of articles. It provides an outline of the text, forming an XML structure that allows for restoring the original PDF layout. Additionally, Grobid enables precise localization of images, tables, formulas, and other data types. With the provided bounding boxes, we can obtain the corresponding images using the PyMuPDF tool 202020https://github.com/pymupdf/PyMuPDF. Further leveraging the OCR recognition integrated into DeepShovel [46], we can convert tables, formulas, and other elements into Markdown format. The parsing process is completed by writing all the parsed content into a markdown file for further use. Throughout the entire process, for tables, we utilize the DeepShovel PDF Table Parser 212121https://github.com/ShaoZhang0115/Table-Extraction-for-Geoscience-Literature. This tool ensures the completeness and accuracy of the table content while preserving the table structure, making it convenient to reconstruct tables using Markdown. As for formulas, we employ an improved version of Latex-OCR 222222https://github.com/lukas-blecher/LaTeX-OCR for the recognition, converting the parsing results into the string format. We open-source our PDF parsing solution on GitHub 232323https://github.com/Acemap/pdf_parser.
@@ 4.2.1 Domain General Natural Language Instruction
For the general instruction learning data, we have integrated four platforms constructed by Acemap, and reconstructed the data accordingly.
@@
This two platforms can be understood as collections of papers and datasets. Therefore, the Related Paper (with abstract) and Reference resolution of Deep Literature, as well as the Reference resolution of DataExpo, serve as excellent datasets for establishing referential relationships.
Using the text processing tool mentioned earlier, we explicitly employ a multi-threaded Grobid to process all documents and convert them into an intermediate XML format. Within the converted XML, we identify the bibl_id of in-text citations and then locate the corresponding reference paper titles in the XML’s reference section.
@@
For the article’s title, we used the Wikipedia library in Python 242424https://github.com/goldsmith/Wikipedia, which supports accessing sections of a Wikipedia page. Each section’s title, text, and sub-sections form a nested structure. By recursively traversing each page, we obtained a list of triplets comprising each section’s level, title, and paragraph. The triplets are structured as (level, title, paragraph), where the level indicates the depth of nesting, the title represents the section’s title, and the paragraph contains the corresponding text content.
@@
To extract the Entity mentioned in the article, we use the requests library and the BeautifulSoup 252525https://www.crummy.com/software/BeautifulSoup/ library to scrape the Wikipedia page directly. We retrieve the text from all tags labeled "p" and "ul" and treat them as paragraphs. Next, within these paragraph tags, we search for tags labeled "a" with a href attribute starting with "/wiki/". These represent the highlighted blue hyperlinked sections within the text. We collect these entities and output the tuple (paragraph, entities).
@@ 4.2.3 Self-Instruct
According to Alpaca [51] and Baize [52], using problem seeds to generate answer from ChatGPT 262626In the data curation process and experiments throughout this paper, we use the 2023 March version of ChatGPT and 2023 March version of GPT-4 unless otherwise specified is an appropriate way to build instruction tuning data. In geoscience scenarios, we generate 1000 questions per subject under the geoscience, and we put the problem seeds on GeoGalactica’s Github Repo.
@@
| Scholar | Title (with Abstract) | (abstract; title) | 2,690,569 | |
@@
| DataExpo | Title | (abstract; title) | 216,036 | |
@@ 5.3 Tool Learning
[linecolor=baby_blue, linewidth=2pt, roundcorner=10pt, fontcolor=black, shadow=true, shadowsize=2pt, shadowcolor=light_grey]
Example 1:
Question: "What is the weather in New York 3M years ago?"
@@
Action Input: "New York, Weather, 3M years" -
Example 2
Question: "What is the definition of plate tectonics?"
@@ 6.1.2 MMLU
Fortunately, we came across Galpaca-30B on Hugging Face 272727https://huggingface.co/GeorgiaTechResearchInstitute/galpaca-30b, which significantly reduced the carbon emissions from our finetuning experiments. This model utilized Alpaca’s instructions to learn from the dataset and was applied to SFT on Galactica-30B. Upon horizontal comparison, Galpaca-30B performed notably worse than the original Galactica and GeoGalactica in the majority of disciplines. This indicates that instruction learning in the general domain can significantly impact the performance of specialized domain models during practical evaluations.